A rocket is moving at a speed of
200m/s. Towards a stationary target while moving it emitts a wave of frequence 1000hz some of sound reaching the target reflected back as an echo calculate 1 the frequency of the sound detected by the target
1
Expert's answer
2018-01-28T09:42:08-0500
The Doppler shift of a frequency
f=(1+v/c) f_0,
where c=340 m/s is a sound speed.
Thus the frequency of the sound detected by the target
f=(1+200/340)1000=1588 Hz